# Particle Acceleration in Collisionless Magnetically Arrested Disks

### v1

@@ Acknowledgements.
Figure 8: Density-averaged energy per particle $\langle\xi\rangle=\sum_{s}\langle\gamma_{s}\rangle\rho_{s}c^{2}/\sum_{s}\rho_{s}$ for particle species $s\in\{e^{-},e^{+},i\}$ shown in panels (a) and (b). These figures correspond to the right sides of the inset panels shown in Fig. 2 of the main text. The green contours correspond to inner light surfaces for the magnetic field orbital velocity $\Omega_{F}=-E_{\theta}/\sqrt{h}B^{r}$, where $h$ is the determinant of the spatial part of the metric. The base inner light surface is rapidly varying in the equatorial region, so to show the global characteristic we display a contour acquired from a smoothed light surface function. The ergosphere is denoted with the dashed magenta contours. The bottom panels (c) and (d) show the amount of unscreened electric field by means of quantity $\bm{D}\cdot\bm{B}/B^{2}$, where $\bm{D}$ and $\bm{B}$ are the FIDO-measured electric and magnetic fields. The colors are log-scaled and simultaneously conserve the sign of the quantity.

### v2

@@ Acknowledgements.
Figure 8: Two distinct particle acceleration channels during a flux eruption (left) showing most activity around the poles, which is followed by the recovering of the flow (right) that encounters an equatorial largely evacuated (bubble) feature. The density-averaged energy per particle $\langle\xi\rangle=\sum_{s}\langle\gamma_{s}\rangle\rho_{s}c^{2}/\sum_{s}\rho_{s}c^{2}$ for particle species $s\in\{e^{-},e^{+},i\}$ shown in panels (a) and (b). These figures correspond to the right sides of the inset panels shown in Fig. 2 of the main text. The green contours correspond to the inner light surface that is determined by the magnetic field orbital velocity $\Omega_{F}=-E_{\theta}/\sqrt{h}B^{r}$, where $h$ is the determinant of the spatial part of the metric. The inner light surface is rapidly varying in the equatorial region, so to show the global characteristic we display a contour acquired from a smoothed light surface function. The ergosphere is denoted with the dashed blue contours. The bottom panels (c) and (d) show the amount of unscreened electric field by means of quantity $\bm{D}\cdot\bm{B}/B^{2}$, where $\bm{D}$ and $\bm{B}$ are the FIDO-measured electric and magnetic fields. The colors are log-scaled and simultaneously conserve the sign of the quantity.